OpenStudy (anonymous):

Simplify each term, if necessary, and then use the distributive property to combine terms, if possible.

OpenStudy (anonymous):

\[\sqrt{75}+4\sqrt{3}\]

OpenStudy (the_fizicx99):

X_X √75 = √25 * √3 √75 = 5*√3 = 5√3 5√3 + 4√3 = 9√3

OpenStudy (anonymous):

Omg, thank you ! :)

OpenStudy (the_fizicx99):

yw